# A critical assessment of the role of topology on protein thermal stability

### v1

@@ 3.2 Protein YbeA and its control system
Results reported in Figure 3 (A-D) show that YbeA folds cooperatively in a two-state manner, while the control CP-YbeA folds non-cooperatively. This lack of cooperativity is particularly evident from the rather gentle slope of the internal energy curve (Figure 3 A), and from the broadness of the heat capacity curve (Figure 3 D). In addition, the melting temperature of CP-YbeA is significantly (15%) lower than that of YbeA. We have also measured the dependence of the free energy on the fraction of native contacts () at the melting temperature for the knotted-unknotted pair. We find that the behavior of the folding transition changes from two-state (for YbeA) to downhill (or barrierless) (for CP-YbeA) (Figure 4 A). These findings, in particular the decreased thermal stability of CP-YbeA, cannot be attributed solely to its unknotted topological state. Indeed, untying the knot in YbeA introduced drastic structural changes and a significant loss of native interactions, which collectively account for the observed behavior.
@@ 3.3 Protein YibK and its control systems
To isolate the role of the topological state on thermal equilibrium properties, particularly thermal stability, we now examine YibK and its control systems. In this case, the two-state nature of the folding transition is conserved across the three model systems (Figure 3E-H). A minor, yet non-negligible difference is observed between YibK and its in silico control system, IS-YibK. In particular, for the latter, the melting temperature is 7% lower than that of YibK (Figure 3H). Since the number of native contacts is not conserved in the unknotted protein, we cannot ascribe its lower thermal stability to its unknotted topological state. In fact, when one compares YibK and CP-YibK, which are nearly structurally identical and share almost the same number of native contacts, it is found that their thermodynamic behavior is practically coincident, with the thermal stability differing by only 1%. In line with these findings, the similarity between the free energy profiles (Figure 4B) is also significantly higher for YibK and its unknotted circular permutant.
@@ 3.4 Comparison with experimental data
The results reported in the previous section show that as the similarity between the knotted protein and its unknotted control increases, the differences in the thermodynamic equilibrium behavior tend to disappear. In particular, they indicate that if one could construct an unknotted control system conserving the same native geometry (i.e., the atomic positions and, consequently, distances between atoms), one should observe the same thermodynamic behavior. These findings support the view that the protein’s topological state does not enhance thermal stability or affect thermodynamic equilibrium properties. From a physics standpoint, this result is expected since the protein’s energy is not a function of the topological state, and, consequently, the equilibrium density of states should be neither, provided the exploration of the conformational space is ergodic. An identical density of states is expected for both knotted and unknotted control systems, which translates into identical thermodynamic behavior.
@@
However, the experimental measurements carried out for YibK and its unknotted control system contradict this argument, indicating that the knotted topology enhances the protein’s thermal stability. Indeed, a decrease in the melting temperature of 22% (measured with differential scanning calorimetry) or 27% (measured with far-UV Circular Dichroism (CD) signals) was reported Danny2019a for the unknotted protein.
Part of the decrease in thermal stability observed experimentally may be due to the fact that the dimer interface for each chain in the dimeric version of CP-YibK is reduced by 13%, which leads to a loss of intermolecular hydrogen bonds and salt bridges near the dimer interface Danny2019a. Since experimental data refers to the dimeric form of the protein, and simulations focused on the monomer, a direct comparison between experimental and computational data is not possible. That said, it is clear that the experimental and computational results offer opposite views on the role of topology in setting YibK’s thermal stability.
A few years ago, Jennings and co-workers reported two distinct timescales for the unfolding and the untying of protein MTTTm (PDB ID: 106d) – a protein of the same family as YibK that features a similar chain length. Their findings showed that the transition to untied states required at least 6 months to occur Jennings2016, while the transition to unfolded states occurs in 2 weeks. This slow unknotting process suggests that, for deeply knotted proteins like YibK, the full unfolding–untying transition may not be observed within the timescale of a DSC experiment, indicating a deviation from equilibrium thermodynamics.
These experimental observations align with insights from computational simulations. As previously noted, it is particularly challenging to achieve equilibrium distributions for complex polymeric structures — such as deeply knotted proteins — using MC simulations based on a move set that preserves the linear topology of the chain, and starting from unfolded and unknotted conformations. For example, in the case of protein Rds3p, which is 51 amino acids shorter than YibK, simulations preserving linear topology required at least an order of magnitude more MC steps to reach equilibrium compared to simulations using move sets that allow the chain to cross itself Especial2022. Notably, for unknotted proteins (e.g., Fn-III and beta-2-microglobulin), the computational cost is independent of the move set Especial2022.
Given that real protein dynamics also preserve the linear topology of the polypeptide chain, it is plausible that the scan rate (i.e., heating rate) used in the DSC experiment was too fast for the system to equilibrate at each temperature, and even more so to allow the population of unknotted conformations. This supports the view that the experimental measurements likely reflect a non-equilibrium distribution lacking unfolded and unknotted conformations, and, consequently, the measured melting temperature represents an upper bound for the real melting temperature.
@@ 4 Conclusions
Over the past 20 years, a remarkably large body of theoretical and experimental work has been dedicated to the study of knotted proteins. A major line of research has been devoted to understanding whether knots may confer any functional advantage to their carriers Jackson2020. Answering this question will provide insight into the evolutionary significance of knotted proteins and may have a potential impact on designing drugs for technological or medical applications.
A major challenge in studying the role of knots in proteins is the construction of a suitable control system, a version of the protein that is structurally similar to the knotted variant but lacks the knot. In the context of lattice models, it is quite straightforward to untie a model protein while introducing minimal changes to the system; previous simulation efforts in the lattice context have shown that a knotted topology does not enhance the carrier’s thermal stability Soler2013. Unfortunately, in real-world proteins, this is not as simple. Here, we employed a simple off-lattice model to show that as the structural similarity between the protein and its control system increases, the difference in thermodynamic equilibrium properties vanishes. In particular, we have provided compelling evidence that the knot in protein YibK does not enhance its thermal stability. This result is in contradiction with recent experimental data on this protein, which reported a moderate increase in thermal stability for the knotted version of Yibk Danny2019a.
Considering that the timescale required for untying a protein knot can be orders of magnitude longer than that for unfolding Jennings2016, and that Monte Carlo simulations face significant challenges in reaching equilibrium distributions for deeply knotted proteins like YibK when starting from unknotted conformations, we propose that the experimental data may reflect an unfolding–folding transition occurring under non-equilibrium conditions.
@@ 6 Declaration of interests
Work supported by UID/04046/2025 Instituto de Biosistemas e Ciências Integrativas Centre Grant from Instituto de Biosistemas e Ciências Integrativas Centre grant from FCT, Portugal. J.N.C.E. acknowledges financial support from FCT, Portugal, through PhD grant SFRH/BD/144345/2019. B.P.T. acknowledges financial support from Fundação Calouste Gulbenkian through the Gulbenkian Novos Talentos Fellowship. This study was also supported by the European Union (TWIN2PIPSA - Twinning for Excellence in Biophysics of Protein Interactions and Self-Assembly, GA 101079147). A part of this work was performed on the computational resources of INCD (http://www.incd.pt) funded by FCT and UE under project LISBOA-01-0145-FEDER-022153. Access was granted by FCT through project 2022.26279.CPCA.A0.

### v2

@@ 3.2 Protein YbeA and its control system
Results reported in Figure 3A–D show that YbeA folds cooperatively in a two-state manner, while the control CP-YbeA folds non-cooperatively. This lack of cooperativity is particularly evident from the rather gentle slope of the internal energy curve (Figure 3A), and from the broadness of the heat capacity curve (Figure 3D). In addition, the melting temperature of CP-YbeA is significantly (15%) lower than that of YbeA. We have also measured the dependence of the free energy on the fraction of native contacts () at the melting temperature for the knotted-unknotted pair. We find that the behavior of the folding transition changes from two-state (for YbeA) to downhill (or barrierless) (for CP-YbeA) (Figure 4A). These findings, in particular the decreased thermal stability of CP-YbeA, cannot be attributed solely to its unknotted topological state. Indeed, untying the knot in YbeA introduced drastic structural changes and a significant loss of native interactions, which collectively account for the observed behavior.
@@ 3.3 Protein YibK and its control systems
To isolate the role of the topological state on thermal equilibrium properties, particularly thermal stability, we now examine YibK and its control systems. In this case, the two-state nature of the folding transition is conserved across the three systems (Figure 3E-H). A minor, yet non-negligible difference is observed between YibK and its in silico control system, IS-YibK. In particular, for the latter, the melting temperature is 7% lower than that of YibK (Figure 3H). Since the number of native contacts is not conserved in the unknotted protein, we cannot ascribe its lower thermal stability to its unknotted topological state. In fact, when one compares YibK and CP-YibK, which are nearly structurally identical and share almost the same number of native contacts, their thermodynamic behavior is practically identical, with the thermal stability differing by only 1%. In line with these findings, the similarity between the free energy profiles (Figure 4B) is also significantly higher for YibK and its unknotted circular permutant.
@@ 3.4 Comparison with experimental data
The results reported in the previous section show that as the similarity between the structures of a knotted protein and its unknotted control increases, differences in thermodynamic equilibrium behavior diminish. Indeed, simulations indicate that an unknotted control preserving the same native geometry (and, consequently, distances between atoms) would exhibit similar thermodynamic behavior. From a physics perspective, this is expected: in our model, the Hamiltonian depends only on the native contact map (atomic positions and corresponding interactions) and is independent of the topological state. All native interactions are identical between the knotted and unknotted structures, and non-native interactions are neutral by construction. Therefore, if the native contact maps of knotted and unknotted structures are identical, (or nearly identical as in the case of YibK and its control system CP-YibK) an identical (or nearly identical) density of states is expected, and, consequently an identical (or nearly identical) thermodynamic behavior as well.
@@
However, the experimental measurements carried out for YibK and CP-YibK do not support the findings from simulations, indicating that the knotted topology enhances the protein’s thermal stability. Indeed, a decrease in the melting temperature of 22%, measured with differential scanning calorimetry (DSC), or 27%, measured with far-UV circular dichroism (CD) signals, was reported 6 for the unknotted protein.
Part of the decrease in thermal stability observed experimentally may be due to a 13% reduction in the dimer interface per chain in the dimeric version of CP-YibK. A fraction of the decreased stability could then be attributed to the weakened dimer interface, featuring fewer intermolecular hydrogen bonds and salt bridges 6. Because experimental measurements were performed on the dimer while simulations focused on the monomer, a direct comparison between experimental and computational data is not straightforward. Furthermore, the thermal unfolding experiments were not demonstrated to be reversible, unlike the simulations. If unfolding is irreversible, the reported experimental melting temperatures reflect apparent rather than true thermal stability. If the control system CP-YibK used in experiments is more prone to aggregation than the knotted YibK, this could also easily suppress the experimental melting temperature. Although simulations provide a controlled framework to isolate the role of protein topology in thermal stability, they rely on approximations at both the level of protein representation and the Hamiltonian. Ultimately, this means that the simulation results should be interpreted with care rather than providing definitive thermodynamic measurements. Despite the limitations inherent in both experimental and computational approaches, their results provide different views of the role of topology in determining YibK’s thermal stability. In the following, we discuss a possible, fundamental reason for this discrepancy.
A few years ago, Jennings et al. 3 reported the existence of two distinct timescales for unfolding and untying, upon chemical denaturation, in the deeply knotted protein MTTTm (PDB ID: 106d), which belongs to the same family as YibK and has a comparable chain length. They found that the transition to the untied states required at least six months, whereas unfolding occurred on a timescale of approximately two weeks. A previous study by Jackson and co-workers reported that the knotted topology is retained in the unfolded state of YibK, following chemical denaturation 30. These experimental observations are consistent with insights from computational simulations based on lattice models that also reported the occurrence of knotted conformations in thermally denatured states of deeply knotted lattice proteins.40.
As discussed in the Model and Methods section, achieving equilibrium distributions for complex polymeric systems, such as deeply knotted proteins, is particularly challenging when using off-lattice Monte Carlo (MC) simulations with a move set that preserves the chain’s linear topology, starting from unfolded, unknotted conformations. For instance, in simulations of the deeply knotted protein Rds3p, which is 51 amino acids shorter than YibK, the linear topology-preserving move set required an order of magnitude more MC steps to reach equilibrium compared to simulations that allow chain crossings 9. By contrast, for unknotted proteins, such as Fn-III and 2-microglobulin, the computational cost is independent of the move set employed 9. The fact that, for unknotted proteins, the equilibration time is independent of the move set, whereas deeply knotted proteins exhibit a strong dependence, indicates that knotting introduces a significant degree of topological frustration, which delays the equilibration time, when chain crossings are forbidden. We repeated the analysis of the equilibration time for YibK and found that enforcing linear topology preservation increases the equilibration time by 1.78 orders of magnitude relative to simulations that do not preserve the linear topology of the chain (Figure 5). Since real protein dynamics also preserves the linear topology of the polypeptide chain, it is plausible that the scan rate (i.e., heating rate) used in the DSC experiments is too fast for the knotted protein to equilibrate at each temperature, and even more so to allow the population of denatured conformations that are also unknotted. Thus, the pronounced separation of timescales for unknotting and unfolding suggests that, for deeply knotted proteins such as YibK, the complete unfolding–untying transition may not be accessible within the time frame of a DSC experiment. In this case, the experimental measurements likely reflect a non-equilibrium distribution lacking unfolded states that are also unknotted; in that case, the measured melting temperature represents an apparent value rather than a true thermodynamic parameter.
@@ 4 Conclusions
Over the past 20 years, a remarkably large body of theoretical and experimental work has been dedicated to the study of knotted proteins13, 20, 19. A major line of research has focused on whether knots confer any functional advantage to their carriers 21. Answering this question will provide insight into the evolutionary significance of knotted proteins and may have an impact on the design of drugs for technological or medical applications.
A major challenge in studying the role of knots in proteins is the construction of a suitable control system, a version of the protein that is nearly identical to the knotted variant but lacks the knot 25. In the context of lattice models, it is quite straightforward to untie a model protein while introducing minimal changes to the structure 11, and previous simulation efforts in the lattice context have shown that a knotted topology does not enhance the protein’s thermal stability 40. Unfortunately, in more realistic off-lattice models, as well as in real-world proteins, engineering unknotted control systems is not as straightforward.
Here, we conducted extensive Monte Carlo simulations of a simple off-lattice model with energetics governed by a Go potential to show that protein knots can strongly affect the equilibration time and make the equilibrium distribution not accessible in a reasonable timescale. Indeed, by considering the chain’s linear topology as either ‘on’ or ‘off’, our simulations corroborated that knots impose this limitation, with the move set that preserves the linear topology requiring almost two orders of magnitude more Monte Carlo steps to equilibrate. Additionally, by comparing the knotted protein YibK with a nearly identical control system experimentally engineered 6, our results show that, if equilibrium is indeed reached, knotting does not enhance the thermal stability of this deeply knotted protein. This finding contradicts experimental data analysing the same pair of structures, which reported a moderate decrease in thermal stability for the unknotted one 6. Based on results from the simulations conducted here, and previous experimental results 30, 3 and computational data 40, we argue that the experimental enhancement of thermal stability in YibK likely reflects non-equilibrium effects. The persistence of knotted conformations in the denatured ensembles leads to a strong separation of timescales between unfolding (couple of weeks) and unknotting (at least 6 months) 3, and suggests that standard thermal denaturation experiments may not be able to access unfolded and unknotted states, leading to an apparent melting temperatures that does not correspond to a an equilibrium distribution. Overall, our findings indicate that protein knots do not intrinsically enhance thermal stability (or any other thermodynamic equilibrium property), and suggest that research in this topic should focus on non-equilibrium properties instead.
@@ 6 Declaration of interests
Work supported by UID/04046/2025 Instituto de Biosistemas e Ciências Integrativas Centre Grant from Instituto de Biosistemas e Ciências Integrativas Centre grant from FCT, Portugal. J.N.C.E. acknowledges financial support from FCT, Portugal, through PhD grant SFRH/BD/144345/2019. B.P.T. acknowledges financial support from Fundação Calouste Gulbenkian through the Gulbenkian Novos Talentos Fellowship. This study was also partially supported by the European Union (TWIN2PIPSA - Twinning for Excellence in Biophysics of Protein Interactions and Self-Assembly, GA 101079147). A part of this work was performed on the computational resources of INCD (http://www.incd.pt) funded by FCT and UE under project LISBOA-01-0145-FEDER-022153. Access was granted by FCT through project 2022.26279.CPCA.A0. We thank Sophie Jackson for insightful comments on our manuscript.